The population mean weight of a candy is 10 grams with a standard deviation of 2 grams. A sample of 100 candies is taken. What is the probability that the sample mean is at least 10.6 grams?
1
Expert's answer
2020-10-27T20:10:06-0400

Let the original distribution function be F, then we need to perform:

"S=\sum\limits_{i=1}^n F(x_i)\geq1060"

From the Central limit theorem, the sum function will also be a normal distribution with a mathematical expectation of 1000 and a standard deviation of 200. Then apply the formula:

"P(S>10600)=Ф(\infty)-Ф(\frac{1060-1000}{200})=\\=0.5-0.1179=0,3821"

Where Ф is the Laplace function, the values are taken from the table